# RU-AI: A Large Multimodal Dataset for Machine Generated Content Detection

### v1

@@ 4. Conclusion
We propose RU-AI as a large multimodal dataset for machine-generated content detection. Compared with other datasets that contain only single or dual modalities, our approach aligns text, image, and voice data with their machine-generated counterparts. To ensure data diversity, we incorporate three datasets across 18 different object categories, using five generative models for each modality. Experimental results from our proposed baseline show that our unified classification model can effectively detect machine-generated content. However, several challenges remain, such as improving cross-modality detection accuracy, addressing model scalability and availability to general public. We hope making RU-AI dataset publicly available to the public provides direction for future research.
5. Ethics Considerations
The collection of RU-AI dataset is in accordance with institutional review board standards and ACM standards of ethics. Our dataset is created from open-sourced datasets and generative models with respect for the privacy and confidentiality of original sources. We publish our dataset at Zenodo with Digital Object Identifier: 10.5281/zenodo.11406538 to encourage open and collaborative research.

### v2

@@ 4. Conclusion
We propose RU-AI, a comprehensive large-scale multimodal dataset designed for machine-generated content detection. Unlike existing datasets, which are limited to single or dual modalities, RU-AI aligns text, image, and voice data with their corresponding machine-generated counterparts. To ensure data diversity, we integrate three large open-source datasets and employ five generative models for each modality, both with and without data augmentation. Experimental results from SOTA baseline models show that unified classification models struggle to effectively detect machine-generated content on our dataset, especially when augmented with noise. This highlights the persistent challenges in the era of generative AI. By making the RU-AI dataset publicly available, we hope to foster further research in this dynamic and rapidly evolving area.